# Slot State Space Models

### v1

@@ 7.1 Multi-Object Video Prediction
Models. We employ the SlotSSM architecture described in Section 4.3, utilizing the Mamba Block [14] as the SSM implementation for both SlotSSM and baseline models. We use the same number of slots for all layers, and thus only apply the Slot Encoder at the first layer. We compare our model against the following baselines:
@@ 7.2 Long-Context Reasoning
Figure 5: Long-Context Construction and Model Efficiency in the Blinking Color Balls Benchmark. Left: Patchifying the context images to provide long-sequence inputs. Right: Comparison of model inference latency. SlotSSM demonstrates its computational efficiency for long-sequence processing tasks.
Blinking Color Balls Benchmark. The benchmark comprises two variants: the Earliest Color and the Most Frequent Color versions. Both variants consist of epoisodes of image sequence ${\bf x}_{1:N}$, partitioned into context images ${\bf x}_{1:N-1}$ and a target image ${\bf x}_{N}$. For each timestep in ${\bf x}_{1:N-1}$, all bouncing balls are initially colored white. Then, one ball is randomly selected and assigned a non-white color, which is randomly sampled from a set of 5 predefined colors The coloring of balls in the target image ${\bf x}_{N}$ depends on the variant:
@@
To transform the dataset into a long-range task, we patchify each context image into non-overlapping patches and provide the flattened patch sequence as input to the model. We present this process in Figure 5(a). Specifically, each context image is divided into $P\times P$ patches, which are then flattened into a sequence of length $P^{2}$. For a context length of $N-1$, this results in a total input sequence length of $L=(N-1)\times P^{2}$. Consequently, the models must identify and track objects from partial views provided by the patches, while simultaneously remembering and counting the color assignments for each object, making the task highly challenging and non-trivial. The final task is to predict the target image given this long sequential input.
Experimental Setup.
We evaluate models on Earliest Color with $N=6$ and Most Frequent Color with $N\in\{6,11\}$, and we employ patch sizes of $P\in\{4,8,16\}$. The resulting input sequence lengths are $L\in\{80,160,320,640,1280,2560\}$. Note that Most Frequent Color with $N=11$ setting poses a greater challenge because the increased context length leads to more color assignments, requiring the models to exhibit stronger memorization and reasoning capabilities.
Models. We employ the same encoder, decoder, and the SlotSSM architectures as described in Section 7.1. For slot encoding, each image patch is treated as a small image and processed by the transformer slot encoder. The resulting slots from the last time step are then provided to the transformer decoder to predict the full target image. We compare our SlotSSM against several baselines:
Single State SSM, SlotTransformer, and RIM. Additionally, we introduce a novel multi-slot design called SlotRNNs, which shares model weights across the RNNs of each slot and uses self-attention layers between time steps as the slot mixer. SlotRNNs can be viewed as a special case of RIMs, with shared RNN weights and dense state updates instead of sparse state updates. Empirically, we find that SlotRNNs exhibit more stable training and improved performance compared to RIMs. We speculate that this improvement may be attributed to the benefits of weight sharing. For a fair comparison, all slot-based models use 6 slots. We also carefully match the model size as in Section 7.1
@@ A.2 Dataset Design
We provide an illustrative example of the dataset design in Figure 8. Each episode of the dataset contains a context-target pair $({\bf x}_{1:N-1},{\bf x}_{N})$. At each timestep in ${\bf x}_{1:N-1}$, all bouncing balls are first colored white, and then one ball is randomly picked and colored with one of 5 non-white colors. This process is repeated for all context frames, and it is represented in the rows in Figure 8(top). Note that the object picking and coloring are performed independently for each timestep, thus one ball could be selected none or multiple times and colorized with the same or different colors across different timesteps.
@@ A.3 Challenges and Qualitative Comparison
Figure 10 presents a qualitative comparison of the models’ performance on the task. The results reveal a clear categorization of the models based on their capability to address the two levels of challenges. The baseline RIM model successfully predicts the object positions in the target image but struggles with learning the color assignment rules. Consequently, it predicts the color white that generally have the highest appearance probability. Note that the rendered images are based on the argmax of the logits over the color categories. Models such as SlotRNN and Single State SSM demonstrate the ability to learn color assignments, but they make mistakes in some cases. In contrast, SlotSSM and SlotTransformer successfully achieve both accurate position prediction and color assignment.

### v3

@@ 7.1 Multi-Object Video Prediction
Models. We employ the SlotSSM architecture described in Section 4.3, and we adopt the Mamba Block [14] as the SSM implementation for both SlotSSM and baseline models. We use the same number of slots for all layers, and thus only apply the Slot Encoder at the first layer. We compare our model against the following baselines:
@@ 7.2 Long-Context Reasoning
Figure 5: Long-Context Construction and Model Efficiency in the Blinking Color Balls Benchmark. Left: We construct long-sequence inputs by patchifying the context images. Right: Comparison of model inference latency with batch size 6. SlotSSM demonstrates computational efficiency for long-sequence processing tasks.
Blinking Color Balls Benchmark. The benchmark comprises two variants: the Earliest Color and the Most Frequent Color versions. Both variants consist of episodes of image sequence ${\bf x}_{1:T}$, partitioned into context images ${\bf x}_{1:T-1}$ and a target image ${\bf x}_{T}$. For each timestep in ${\bf x}_{1:T-1}$, all bouncing balls are initially colored white. Then, one ball is randomly selected and assigned a non-white color, which is randomly sampled from a set of 5 predefined colors The coloring of balls in the target image ${\bf x}_{T}$ depends on the variant:
@@
To transform the dataset into a long-range reasoning task, we patchify each context image into non-overlapping patches and provide the flattened patch sequence as input to the model. We present this process in Figure 5(a). Specifically, each context image is divided into $P\times P$ patches, which are then flattened into a sequence of length $P^{2}$. For a context length of $T-1$, this results in a total input sequence length of $L=(T-1)\times P^{2}$. Consequently, the models must identify and track objects from partial views provided by the patches, while simultaneously remembering and counting the color assignments for each object, making the task highly challenging and non-trivial. The final task is to predict the target image given this long sequential input.
Experimental Setup. We evaluate models on Earliest Color with $T=6$ and Most Frequent Color with $T\in\{6,11\}$, and we employ patch sizes of $P\in\{4,8,16\}$. The resulting input sequence lengths are $L\in\{80,160,320,640,1280,2560\}$. Note that Most Frequent Color with $T=11$ setting poses a greater challenge because the increased context length leads to more color assignments, requiring the models to exhibit stronger memorization and reasoning capabilities.
Models. We employ the same encoder, decoder, and the SlotSSM architectures as described in Section 7.1. For slot encoding, each image patch is treated as a small image and processed by the transformer slot encoder. The resulting slots from the last time step are then provided to the transformer decoder to predict the full target image. We compare our SlotSSM against several baselines: Single State SSM, SlotTransformer, and RIM. Additionally, we introduce a novel multi-slot design called SlotRNNs, which shares model weights across the RNNs of each slot and uses self-attention layers between time steps as the slot mixer. SlotRNNs can be viewed as a special case of RIMs, with shared RNN weights across states and dense state updates instead of sparse state updates. Empirically, we find that SlotRNNs exhibit more stable training and improved performance compared to RIMs. For a fair comparison, all slot-based models use 6 slots. We also carefully match the model size as in Section 7.1
@@ A.2 Dataset Design
We provide an illustrative example of the dataset design in Figure 8. Each episode of the dataset contains a context-target pair $({\bf x}_{1:T-1},{\bf x}_{T})$. At each timestep in ${\bf x}_{1:T-1}$, all bouncing balls are first colored white, and then one ball is randomly picked and colored with one of 5 non-white colors. This process is repeated for all context frames, and it is represented in the rows in Figure 8(top). Note that the object picking and coloring are performed independently for each timestep, thus one ball could be selected none or multiple times and colorized with the same or different colors across different timesteps.
@@
Finally, as illustrated in Figure 5(a), we transform the conditional image generation task into a long-range reasoning task by using patchified context images as input. Instead of providing the $T-1$ context images directly to the model, we flatten non-overlapping patches of the original images to create a long input sequence. Given $P\times P$ patches per image, the context length becomes $L=(T-1)\times P^{2}$. Note that patchification is used intentionally to construct long sequences for the benchmark; SlotSSMs in general do not inherently require patchified inputs and instead use a Slot Encoder to extract slots as input at each time step.
@@ A.3 Challenges and Qualitative Comparison
Figure 10 presents a qualitative comparison of the models’ performance on the task. The results reveal a clear categorization of the models based on their capability to address the two levels of challenges. The baseline RIM model successfully predicts the object positions in the target image but struggles with learning the color assignment rules. Consequently, it predicts the color white that generally have the highest appearance probability for all objects. Note that the rendered images are based on the argmax of the logits over the color categories. Models such as SlotRNN and Single State SSM demonstrate the ability to learn color assignments, but they make mistakes in some cases. In contrast, SlotSSM and SlotTransformer successfully achieve both accurate position prediction and color assignment.